Waterways Corporation is continuing its budget preparations. Waterways had the following static budget and actual overhead coWATERWAYS CORPORATION Manufacturing Overhead Flexible Budget For the Month of MarchPrepare a flexible budget report showing the differences (favorable and unfavorable) in manufacturing overhead costs for thePrepare a responsibility report for the manufacturing overhead for March, assuming only variable costs are controllable. WATE
Answer #1

A F Computation of cost per unit Computation -5700/114000 - 13680/ 114000 - 11400/114000 -6840/ 114000 1 Rate 2 Particulars 3E Manufacturing Overhead Flexible Budget Report 28 For the month of March 29 Budget Actual 30 115000 Production in units 1150A C E Waterways Corporation Responsibility Report 47 48 Manufacturing overhead 49 For the month of March 50 5,750 S 13,800 SA В E F Computation of cost per unit Computation 5700/114000 - 13680/ 114000 1 Rate =5700/114000 = 13680/114000 = 11400 /1140В C E Manufacturing Overhead Flexible Budget Report 28 For the month of March 29 Budget 115000 Actual 30 115000 Production inA В C E Waterways Corporation 47 Responsibility Report 48 Manufacturing overhead 49 For the month of March 50 5735 +B51-C51 +